# Exploring nuclear force with pulsar glitch observation

### v1

@@ 3.2 Constraint from Vela 2000
Figure 10: Fraction of the coupled vorticity $Y_{\rm{gl}}$ (upper panels) and observed step in spin-down rate $\Delta\dot{\Omega}_{\rm{p}}/\dot{\Omega}_{\rm{p}}$ (lower panels) as a function of NS mass $M/M_{\odot}$ for DD-ME2 (left panels), PKDD(middle panels), and NL3 (right panels) with $L_{0}=$30, 40, 60, 80 MeV and $\beta=$ 2 and 3. The dotted line represent the upper bound of measured value for the Vela 2000 glitch, together with the $1\sigma$ deviation.
@@
Figure 11: (a)Pinning forces $f_{\rm{pin}}$ and (b) observed step in spin-down rate $\Delta\dot{\Omega}_{\rm{p}}/\dot{\Omega}_{\rm{p}}$ for different vortex length. The effective interaction is DD-ME2 with $L_{0}=30$ MeV. The polarization is taken as $\beta=2$.
@@
We assume that the angular velocity lag $\Delta\Omega_{\rm{cr}}$ between the normal component and the superfluid component inside the NSs changes smoothly. A large number of vortices accumulate in the shell near the peak close to the crust-core interface. When these vortices are suddenly unpinned due to $\Delta\Omega_{\rm{cr}}$ exceeding its maximum value, a glitch occurs and $\Delta\Omega_{\rm{cr}}$ decreases. The mechanism means that the vortices near the peak at the outer boundary of the inner crust cannot be unpinned due to their extreme large lag. Therefore, the weak pinning region of an NS have a neglect contribution to the triggering of glitch and its successive relaxation. We further assume that vortices in the region where $R_{\rm{L}}^{*}<r<R_{\rm{M}}^{*}$ have been completely removed by the Magnus force after the glitch is triggered, where $R_{\rm{M}}^{*}$ is the location of the minimum of the $\Delta\Omega_{\rm{cr}}$ profile (between two peaks); then the angular momentum transfer can be immediately evaluated via Eq. (31) and setting $r_{\mathrm{max}}=R_{\rm{L}}^{*}$ and $r_{\mathrm{o}}=R_{\rm{M}}^{*}$.
@@ 3.3 The effect of vortex length
In previous calculations, we have assumed a vortex length of $l=5000R_{\rm{ws}}$. Based on Seveso et al. (2016), we also investigate the impact of $l$ on our results. In Fig. 11(a), we show pinning forces as a function of density for $l=1000R_{\rm{ws}},~{}2500R_{\rm{ws}}$, and $5000R_{\rm{ws}}$. We can see that the pinning force increases as $l$ decreases. For smaller $l$, the pinning force does not show a significant decline near the crust-core density. Fig. 11(b) illustrates the constraints from Vela 2000 for different $l$. Clearly, Vela 2000 does not support a short vortex length, which implies that the vortex tension is strong.
@@ 3.4 The effect of pinning force strength
Since EoSs and pinning forces are self-consistently calculated in this work, Vela 2000 provides a strong constraint on the NS mass. We now temporarily break the relation between EoS and pinning force and discuss the constraints from Vela 2000 under different pinning force strengths. According to Eqs. (31) and (32), when EoS and NS mass are given, the angular momentum transfer $\Delta L$ during a glitch depends only on the distribution of the pinning force within the inner crust. We use the scaling factor $R_{f}$ to parameterize the pinning force, i.e., $f_{\rm{pin}}(R_{f})=R_{f}f_{\rm{pin}}$, so the angular momentum transfer depends only on $R_{f}$.
In Fig. 12, we present the angular momentum transfer $\Delta L$ as a function of NS mass for different $R_{f}$. The effective interaction and the original polarization are taken as DD-ME2 with $L_{0}$ and $\beta=3$, respectively. In Fig. 12, the upper and lower black solid lines correspond to the lower and upper limits of angular momentum transfer; the latter are obtained by setting $Y_{\rm{gl}}=0$ and $Y_{\rm{gl}}=1$ in Eq. (32), respectively. The angular momentum transfers that validate the snowplow model are shaded by gray. By substituting the upper and lower limits of $\Delta\dot{\Omega}_{\rm{p}}/\dot{\Omega}_{\rm{p}}$ into Eqs. (32) and (33), we can obtain the lower and upper limits of angular momentum transfer that are consistent with the observations. The green dashed line in Fig. 12 draw the upper limit of angular momentum transfer derived from the observed lower limit. Then, the angular momentum transfer during a glitch should lie within the green region shown in Fig. 12 in order to simultaneously satisfy the snowplow model requirements and the observations from Vela 2000. We can see that the angular momentum transfer decreases as the NS mass increases. This is because the heavier NSs correspond to the smaller radius, which can reduce the moment of inertia available for angular momentum transfer after the vortex avalanche. The angular momentum transfer decreases with decreasing pinning force , allowing lighter NSs with thicker crusts to match observations of Vela 2000. As shown in Fig. 12, when $R_{f}$ is reduced from 1 to 0.5 and 0.2, Vela 2000 constrains its mass from about 2.3 $M_{\odot}$ to 1.9 $M_{\odot}$ and 1.4 $M_{\odot}$, respectively.
@@ 4 Summary and Perspective
Pulsar glitch study has the potential to not only provide unique constraints on the dense matter EoS, but also shed light on the composition and state of matter at supranuclear densities. In this work, to maximise the information extracted from glitch observations, we have demonstrated an attempt to explore nuclear force in the context when all required ingredients are properly represented. In practice, we refit the effective interactions DD-ME2, PKDD, and NL3 to the nuclear saturation properties, and then they are used to perform the calculations of the unified EoSs, composition, and global properties of NSs in a consistent manner. Using the semi-classical method, the evolutions of associated pinning force acting on the vortex are determined with the dependence to the nuclear parameters and superfluid polarization strength. Taking these microscopic and macroscopic inputs to the snowplow model for pulsar glitch, the constraints to nuclear force and Vela mass from Vela 2000 glitch are conducted.
@@
We find that vortices favor interstitial pinning with weak strength in the top of the inner crust; while vortices strongly pinning to nuclei in the bottom. In the isoscalar channel, the nuclear forces have slight impact to the pinning properties, at least for these forces we choose. In the isovector channel, the smaller symmetry energy slope leads to the weaker pinning for interstitial pinning case and to the stronger pinning for nuclear pinning case. The stronger polarization reduces the pinning strength for both interstitial and nuclear pinning cases. The observation of Vela 2000 glitch constrains the symmetry energy slope to below 40 MeV, and indicates that its own is a massive pulsar.
@@
With advancements in many-body theory and observational instruments, future efforts should focus on reducing model uncertainties and fostering multi-wavelength collaborations in glitch data analysis (Liu et al., 2024a, b). On one hand, non-relativistic and relativistic full quantum calculations, accounting for uncertainties in nuclear forces, can provide more accurate and detailed insights into pinning energies within the inner crust. Such calculations would be more realistic if one incorporates the effects of pasta structures at the base of the inner crust. On the other hand, high-accuracy facilities like FAST, SKA, eXTP, with the help of multi-wavelength collaborative observations will offer a more unified understanding of stellar interiors and magnetospheres. These advancements may lead to the development of a universal glitch model for Vela-like pulsars, offering a comprehensive description of internal vortex dynamics. This synergy between theoretical and observational progress could greatly enhance our understanding of the nuclear force, particularly its poorly understood isospin dependence, as revealed in the present study.
@@ Acknowledgments
The work is supported by National SKA Program of China (No. 2020SKA0120300), National Natural Science Foundation of China (grant Nos. 12273028, 12494572).

### v2

@@ 3.2 Constraints from the 2000 Vela glitch
Figure 10: Fraction of the coupled vorticity $Y_{\rm{gl}}$ (upper panels) and observed step in spin-down rate $\Delta\dot{\Omega}_{\rm{p}}/\dot{\Omega}_{\rm{p}}$ (lower panels) as a function of NS mass $M/M_{\odot}$ for DD-ME2 (left panels), PKDD(middle panels), and NL3 (right panels) with $L_{0}=$30, 40, 60, 80 MeV and $\beta=$ 2.0 and 3.0. The dotted line represent the upper bound of measured value for the 2000 Vela glitch, together with the $1\sigma$ deviation.
@@
Figure 11: (a)Pinning forces $f_{\rm{pin}}$ and (b) observed step in spin-down rate $\Delta\dot{\Omega}_{\rm{p}}/\dot{\Omega}_{\rm{p}}$ for different vortex length. The effective interaction is DD-ME2 with $L_{0}=30$ MeV. The polarization is taken as $\beta=3.0$.
@@
We assume that the angular velocity lag $\Delta\Omega_{\rm{cr}}$ between the normal component and the superfluid component inside the NSs changes smoothly. A large number of vortices accumulate in the shell near the peak close to the crust-core interface. When these vortices are suddenly unpinned due to $\Delta\Omega_{\rm{cr}}$ exceeding its maximum value, a glitch occurs and $\Delta\Omega_{\rm{cr}}$ decreases. The mechanism means that the vortices near the peak at the outer boundary of the inner crust cannot be unpinned due to their extreme large lag. Therefore, the weak pinning region of an NS have a negligible contribution to the triggering of glitch and its successive relaxation. We further assume that vortices in the region where $R_{\rm{L}}^{*}<r<R_{\rm{M}}^{*}$ have been completely removed by the Magnus force after the glitch is triggered, where $R_{\rm{M}}^{*}$ is the location of the minimum of the $\Delta\Omega_{\rm{cr}}$ profile (between two peaks); then the angular momentum transfer can be immediately evaluated via Eq. (31) and by setting $r_{\mathrm{max}}=R_{\rm{L}}^{*}$ and $r_{\mathrm{o}}=R_{\rm{M}}^{*}$.
@@
Haskell et al. (2013) also applied the snowplow model to the 2000 Vela glitch assuming strong pinning between the flux tubes and neutron vortices. Our results can be compared with theirs using the fraction of pinned vorticity in the core set to zero. In Haskell et al. (2013), the results calculated with most EoSs and $Y_{p}$ match the observed value. However, our results are hard to reconcile with the observed data, and fitting them is only possible for several effective interactions with $L_{0}<40$ MeV and strong polarization in a narrow interval of masses. One reason is that the location of the maximum $\Delta\Omega_{\rm{cr}}$ is close to the bottom of the inner crust, as shown in Fig. 9. From Eq. (31), the angular momentum transfer increases when the crust region supported for the angular momentum transfer is expanded. Therefore, compared to Haskell et al. (2013), our calculations indicate that the inner crust contains a larger angular momentum transfer region, leading to more angular momentum transfer, finally resulting in an increase in $Y_{\rm{gl}}$ and a decrease in $\Delta\dot{\Omega}_{\rm{p}}/\dot{\Omega}_{\rm{p}}$.
@@ 3.3 The effect of vortex length
The rigidity length $l$ of vortex is related to its tension. A stronger tension reduces the ability of a vortex to bend and pin to a pinning configuration, resulting in a lower the pinning force. In previous calculations, we have assumed a vortex length of $l=5000R_{\rm{ws}}$. Based on Seveso et al. (2016), we also investigate the impact of $l$ on our results. In Fig. 11(a), we show pinning forces as a function of density for $l=1000R_{\rm{ws}},~{}2500R_{\rm{ws}}$, and $5000R_{\rm{ws}}$. We can see that the pinning force increases as $l$ decreases. For smaller $l$, the pinning force does not show a significant decline near the crust-core density. Fig. 11(b) illustrates the constraints from the 2000 Vela glitch for different $l$. Clearly, Vela 2000 does not support a short vortex length, which implies that the vortex tension is strong.
@@ 3.4 The effect of pinning force strength
Since EoSs and pinning forces are self-consistently calculated in this work, the 2000 Vela glitch provides a strong constraint on the NS mass. We now temporarily break the relation between EoS and pinning force and discuss the constraints from the 2000 Vela glitch under different pinning force strengths. According to Eqs. (31) and (32), when EoS and NS mass are given, the angular momentum transfer $\Delta L$ during a glitch depends only on the distribution of the pinning force within the inner crust. We use the scaling factor $R_{f}$ to parameterize the pinning force, i.e., $f_{\rm{pin}}(R_{f})=R_{f}f_{\rm{pin}}$, so the angular momentum transfer depends only on $R_{f}$.
In Fig. 12, we present the angular momentum transfer $\Delta L$ as a function of NS mass for different $R_{f}$. The effective interaction and the polarization are taken as DD-ME2 with $L_{0}=30$ MeV and $\beta=3.0$, respectively. In Fig. 12, the upper and lower black dashed lines correspond to the lower and upper limits of angular momentum transfer; the latter are obtained by setting $Y_{\rm{gl}}=0$ and $Y_{\rm{gl}}=1$ in Eq. (32), respectively. The angular momentum transfers that validate the snowplow model are shaded by gray. By substituting the upper and lower limits of $\Delta\dot{\Omega}_{\rm{p}}/\dot{\Omega}_{\rm{p}}$ into Eqs. (32) and (33), we can obtain the lower and upper limits of angular momentum transfer that are consistent with the observations. The green dashed line in Fig. 12 draws the upper limit of angular momentum transfer derived from the observed lower limit. Then, the angular momentum transfer during a glitch should lie within the green region shown in Fig. 12 in order to simultaneously satisfy the snowplow model requirements and the observations from the 2000 Vela glitch. We can see that the angular momentum transfer decreases as the NS mass increases. This is because the heavier NSs correspond to the smaller radius, which can reduce the moment of inertia available for angular momentum transfer after the vortex avalanche. The angular momentum transfer decreases with decreasing pinning force , allowing lighter NSs with thicker crusts to match observations of the 2000 Vela glitch. As shown in Fig. 12, when $R_{f}$ is reduced from 1 to 0.5 and 0.2, the 2000 Vela glitch constrains its mass from about 2.3 $M_{\odot}$ to 1.9 $M_{\odot}$ and 1.4 $M_{\odot}$, respectively.
@@ 4 Summary and Perspective
Pulsar glitch study has the potential to not only provide unique constraints on the dense matter EoS, but also shed light on the composition and state of matter at supranuclear densities. In this work, to maximise the information extracted from glitch observations, we have demonstrated an attempt to explore nuclear force in the context when all required ingredients are properly represented. In practice, we refit the effective interactions DD-ME2, PKDD, and NL3 to the nuclear saturation properties, and then they are used to perform the calculations of the unified EoSs, composition, and global properties of NSs in a consistent manner. Using the semi-classical method, the evolutions of associated pinning force acting on the vortex are determined with the dependence to the nuclear parameters and superfluid polarization strength. Taking these microscopic and macroscopic inputs to the snowplow model for pulsar glitch, the constraints to nuclear force and Vela mass from the 2000 Vela glitch are conducted.
@@
We find that vortices favor interstitial pinning with weak strength in the top of the inner crust; while vortices strongly pinning to nuclei in the bottom. In the isoscalar channel, the nuclear forces have slight impact to the pinning properties, at least for these forces we choose. In the isovector channel, the smaller symmetry energy slope leads to the weaker pinning for interstitial pinning case and to the stronger pinning for nuclear pinning case. The stronger polarization reduces the pinning strength for both interstitial and nuclear pinning cases. The observation of the 2000 Vela glitch constrains the symmetry energy slope to below 40 MeV, and indicates that it is a pulsar with a large mass.
@@
In future work, we will further incorporate entrainment effects (Chamel, 2013, 2024) and fit the improved results to the phenomena at the early-stage of the 2016 Vela glitch (Ashton et al., 2019; Pizzochero, P. M. et al., 2020), such as the glitch rise time and the rapid recovery of overshooting. This aims to place stringent constraints on the glitch model and nuclear forces. Meanwhile, with advancements in many-body theory and observational instruments, future efforts should focus on reducing model uncertainties and fostering multi-wavelength collaborations in glitch data analysis (Liu et al., 2024a, b). On one hand, non-relativistic and relativistic full quantum calculations, accounting for uncertainties in nuclear forces, can provide more accurate and detailed insights into pinning energies within the inner crust. Such calculations would be more realistic if one incorporates the effects of pasta structures at the base of the inner crust. On the other hand, high-accuracy facilities like FAST, SKA, eXTP, with the help of multi-wavelength collaborative observations will offer a more unified understanding of stellar interiors and magnetospheres. These advancements may lead to the development of a universal glitch model for Vela-like pulsars, offering a comprehensive description of internal superfluid dynamics. This synergy between theoretical and observational progress could greatly enhance our understanding of the nuclear force, particularly its poorly understood isospin dependence, as revealed in the present study.
@@ Acknowledgments
We are thankful to K. Sekizawa and the XMU neutron star group for helpful discussions. The work is supported by National SKA Program of China (No. 2020SKA0120300), National Natural Science Foundation of China (grant Nos. 12273028, 12494572).